Two point charges repels each other with a force F when placed in water of dielectric constant 81.what will the force between them when placed the same distance apart in air

Grade:12

2 Answers

Aamina
7 years ago
If some dielectric is present in the space between the charges, the net force acting on each charge is altered becuse charges are induced in the molecules of the intervening medium the force decreases K times.
                        
                                  F=   1    q1q                                           (in vacuum )
                                       4\tiny \prod\small \varepsilon0  r2                                           
 
                              F=      1     . q1q2                                         (in medium)
                                     4\tiny \prod\varepsilon0k     r2              
 
K is dielectric constant.
Hope this helps you.
                                 
Abhinav Singh
6 years ago
Dielectric constant of water.k=81F(air)=1/4pie epsilon×q1q2/r2(square)F (water)=1/4pie epsilon k (dielectric constant)×q1q2/r2 (square)F (water)=1/81×1/4pie epsilon ×q1q2/r2 (square)F (water)=1/81×F (air)Therefore:F (air)=81 F (water).ans.